OpenStudy (anonymous):

The quotient of a number and four is twelve less than the number. What is the number?

OpenStudy (anonymous):

n / 4 = n - 12 16 / 4 = 4 16 - 12 = 4 n = 16

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

To order tickets for a show, there is a $5.00 service fee, plus the cost of the tickets. If the cost of the whole order is $105.00, and 8 tickets were ordered, how much does each ticket cost?